# Text-to-SQL Task-oriented Dialogue Ontology Construction

### v1

@@ 4.7 TOD Ontology Construction Qualitative Analysis
In the second example, we illustrate the impact of adding the DST step, which clarifies the distinction between new dialogue information and existing DB content. Without DST, the model inserts a new entry for “Alexander Bed and Breakfast”. With DST, it updates the existing DB entry by adding only the missing information.
@@
Column value examples help the model generate more general intents and actions, e.g., find_pool” instead of “find_swimming_in_east”. Mentioning dialogue success encourages system actions like “ask_clarification” that support user goals. See Section˜A.2 for predicted ontology excerpts from the “restaurant” (MultiWOZ) and “flights” (SGD) tables.
@@ 4.8 Discussion and Future Work
Our results show that SQL’s structured format improves ontology quality when LLMs interact with a database. The SQL-enhanced model aligns values more accurately using semantic matching or column value examples. Dialogue theory distinguishes existing database information from new input, making the database more user-focused. TeQoDO significantly outperforms the supervised fine-tuned DORE and GenDSI. It also reduces sensitivity to dialogue order. Future work will enable database restructuring through a global view to minimise order effects, since the model could unify table names instead of clustering, as done on ArXiv. Smaller models could improve scalability.
Regarding training data contamination, the results of the direct update approach—with its large variance in table naming—suggest that the model has not memorised the TOD datasets, as it cannot recall the exact table names for each dialogue
For larger or structurally different ontologies than TOD, our results show TeQoDO can generalise, though SQL struggles to express multiple hierarchy levels. We aim to scale TeQoDO to diverse ontology structures by adapting the text-to-SQL component; however, its sequential nature limits parallelisation. The large batch size for concatenated Wikipedia articles may affect performance, though evaluating this is computationally expensive. Finally, our approach is sensitive to prompt phrasing, a common issue with large language models Razavi et al. (2025).
Our evaluation captures all necessary information and is more fine-grained than existing methods by considering the hierarchy levels of TOD ontologies: domains, slots, values, system actions, and user intents. Using an off-the-shelf similarity model with a fixed threshold may introduce bias, though qualitative analysis confirms alignment with ontology quality. In future work, we aim to incorporate human judgement into evaluation to reduce reliance on fixed thresholds.
Dennett (1987) distinguishes competence from comprehension, arguing that the latter is not required for the former. Inspired by this, we view LLMs’ ability to perform tasks like SQL generation as a form of competence to extract task knowledge in the form of an ontology.
We see this work as an important step in distilling human-readable knowledge from otherwise black-box LLMs. Extracting ontologies automatically with LLMs and analysing their structure may enhance their interpretability in downstream tasks. For instance, we might hope to detect hallucinations based on the constructed ontology.
@@ 5 Conclusion
We introduce TeQoDO, a method that uses large language models to build ontologies from task-oriented data, exploiting the inherent SQL programming abilities of LLMs and incorporating dialogue modelling theory. We evaluate TeQoDO on two widely used TOD datasets and find that each proposed step improves performance, surpassing current SOTA. We also show that TeQoDO generalises to large ontologies with different structures. Our results motivate further exploration of this approach for explainability.
@@ A.1 TeQoDO Prompt
    Identify which tables from {db_result_input} are relevant to the new dialogue(s) based on: User intents expressed in the dialogue(s). System actions performed in response. Specific information or entities being queried or discussed. For each relevant table, generate the following SQLite command to inspect its schema: PRAGMA table_info(<table_name>); This will allow you to understand the structure (columns and data types) of the tables you will be working with. Do not create or modify tables yet — only inspect existing ones using PRAGMA.
@@
Baseline: ${}\mathcal{C}_{i}=L(d_{i}),\quad\forall d_{i}\in\mathcal{D},\quad\mathcal{O}_{i}=\mathcal{C}_{0}\cup\dots\cup\mathcal{C}_{i},\quad\mathcal{O}_{0}=\{\varnothing\}$
Iterative Baseline: ${}\mathcal{C}_{i}=L(d_{i},\mathcal{O}_{i-1}),\quad\mathcal{O}_{i}=\mathcal{O}_{i-1}\cup\mathcal{C}_{i},\quad\mathcal{O}_{0}=\{\varnothing\}$
Tracking: $\mathcal{C}_{i}=L(d_{i},\mathbf{b}_{i}),\quad\mathbf{b}_{i}=$ $\mathcal{O}_{i-1}(d_{i})=L(d_{i},\mathcal{O}_{i-1},L(d_{i},\mathcal{O}_{i-1}))$ , ${}\quad\mathcal{O}_{i}=\mathcal{O}_{i-1}\cup\mathcal{C}_{i},\quad\mathcal{O}_{0}=\{\varnothing\}$
Success: $\mathcal{C}_{i}=L(d_{i},\mathbf{b}_{i})$,  ${}\mathbf{b}_{i}=\mathcal{O}_{i-1}(d_{i},\text{is\_successful}(d_{i})),\quad\mathcal{O}_{i}=\mathcal{O}_{i-1}\cup\mathcal{C}_{i},\quad\mathcal{O}_{0}=\{\varnothing\}$
@@ A.2 Predicted Ontology Excerpts
Figure 5: JSON representation of restaurant table predicted via TeQoDO on MultiWOZ test set.
@@
Figure 6: JSON representation of restaurant table predicted via TeQoDO on SGD test set.
@@ A.3 Example Dialogues
See Figure˜7 for the example input dialogues from MultiWOZ test set for the generated queries in Table˜6.
@@ B.1 TOD Ontology Results per Nodeclass
B.2 Prompt as Hyperparameter: Search Plots
See Figures˜8 and 9 for differences in performance for ChatGPT explanation-based prompts and success position in the prompts.
Figure 8: Continuous F1 Performance of different Prompts on MultiWOZ test-set based on LLM explanation.
Figure 9: Continuous F1 Performance of Prompts on MultiWOZ test-set with success mentioned in different steps.

### v2

@@ 4.8 TOD Ontology Qualitative Analysis
See Table˜6 for a qualitative comparison of update queries on MultiWOZ test dialogues, detailed in Section˜A.3. In the first example, we illustrate the impact of adding the DST step, which clarifies the distinction between new dialogue information and existing DB content. Without DST, the model inserts a new entry for “Alexander Bed and Breakfast”. With DST, it updates the existing DB entry by adding only the missing information.
In the second example, “Allenbell” appears in the dialogue, while “The Allenbell” is stored from a previous dialogue. Without similarity matching, the model queries “Allenbell”, yielding no result due to the missing “The”. Similarity matching retrieves “The Allenbell” from the DB. The model chooses which results to include, as less related concepts may also be retrieved.
@@
Column value examples help with more general intents and actions, e.g., find_pool” instead of “find_swimming_in_east”. Mentioning dialogue success encourages system actions like “ask_clarification” that support user goals. See Section˜A.2 for predicted ontology excerpts.
To estimate the frequency of the observations, we sample $50$ dialogues from the MultiWOZ test set and relate qualitative observations to quantitative improvement. For example, if a qualitative effect on user intents appears in $2/50$ dialogues, we expect the corresponding metric to improve by roughly $4\%$.
In $9/50$ cases $(18\%)$, query updates + DST updates rather than inserting, matching the macro F1 rising from $50.1$ to $58.8$, $15\%$ relative. In $2/50$ cases $(4\%)$, query updates with DST and similarity matching yield an update based on the expanded DB results. This better aligns user intents and system actions, leading to a $3\%$ macro gain over DST alone. In $8/50$ cases $(16\%)$, adding DST and column-value examples produces more general user-intent names. This corresponds to a relative macro F1 gain of about $20\%$. Over DST alone, we observe a $4\%$ gain. In $3/50$ cases $(6\%)$, additionally adding success produces more user-focussed system actions. Here, intent F1 improves by $20\%$ and action F1 by around $10\%$ compared to DST plus examples. This shows that the observed behaviour is in line with the ablation.
4.9 Open-source Model Inference
To show the generalisation of TeQoDO to other LLMs, we use it with four open-source models of varying sizes using Vertex AI [^7] https://cloud.google.com/vertex-ai. We use LLaMa-4-scout MetaAI2025_Llama4Blog, a multimodal Mixture of Experts (MoE; Jacobs1991AdaptiveMixtures; Shazeer2017OutrageouslyLargeMoE) with 109B total and 17B active parameters per forward pass. We also use Qwen-3-235B qwen3technicalreport with 22B active parameters and GPT-OSS-20B OpenAI2025_GPT-OSS with 3.6B active parameters. As a small non-MoE model, we use LLaMa-3.1-8B MetaAI2024_LLaMA31Blog.
See Table˜7 for the results, where we additionally show the exact match text-to-SQL performance in multi-turn instructions, based on laban2025llms sharded instruction set-up on the Spider text-to-SQL dataset yu-etal-2018-spider. All models generate useful SQL statements to build a database from the dialogues, showing TeQoDO’s usability across LLMs. The larger models, LLaMa-4-scout and Qwen-3, perform better with the query and update pipeline. Including the DST step and success further enhances performance, approaching GPT-4o-mini on SGD, despite being less capable in text-to-SQL.
Interestingly, the improvement of LLaMa models is less pronounced on MultiWOZ. Here, LLaMa models frequently generate CREATE TABLE queries even with the query-update pipeline. The small LLaMa-3.1-8B model generates CREATE TABLE queries without IF NOT EXISTS, causing overprediction of tables and lower performance than other models. However, on SGD, LLaMa-3.1-8B improves with TeQoDO, as the generated table names have less variation.
For GPT-OSS-20B, which has the fewest active parameters, the DST step and dialogue success in the full TeQoDO pipeline do not improve performance. We assume the small expert cannot model state and dialogue success while interacting with the SQL database. This is reflected in lower domain and user intent performance when using the DST step and dialogue success. Qualitatively, the model overengineers the database, generating more update queries and more specific domains and user intents, which decreases performance.
These results demonstrate the generalisability of TeQoDO to open models as small as 8B parameters and an MoE model with 3.6B active parameters. Moreover, while some text-to-SQL expertise of the LLM is necessary for this approach to work, TeQoDO does compensate to a certain extent for LLMs that lack strong text-to-SQL abilities. The query-update pipeline and the use of DST and success benefit larger models more.
4.10 Prompt phrasing Analysis
Figure 4: Continuous F1 of different Prompts on MultiWOZ test-set based on LLM explanation.
We begin with a manually written prompt following OpenAI best practices (Footnote˜6). Next, we prompt GPT-4o to explain its understanding of this initial prompt. Finally, the LLM revises the prompt using its explanation, facilitating task adherence for LLMs, similar to Kim2024ReEx.
The explanation-revised prompt in Figure˜4 improves MultiWOZ performance by around 10% over the manual prompt. This gain comes from increased clarity – shorter sentences, bullet points, and direct language – which yields more general database updates. For instance, the revised prompt clearly separates desired and non-desired behaviours. The explanation alone performs worse, as it lacks explicit instructions.
Mentioning success only at the end of the update prompt improves performance over not mentioning it, particularly for intents and actions. If mentioned too early, the model ignores the user goal in the update. This shows that performance depends on the phrasing of the prompt, a general issue for LLMs (razavi2025promptsensitivity), but TeQoDO still works with different prompts.
@@ 4.11 Discussion and Future Work
Our results show that SQL’s structured format improves ontology quality when a variety of LLMs with different levels of text-to-SQL abilities interact with a DB. The SQL-enhanced model aligns values more accurately using semantic matching or column value examples. Modular TOD system concepts help in distinguishing existing database information from new input. TeQoDO significantly outperforms the supervisedly fine-tuned DORE and GenDSI. It also reduces sensitivity to dialogue order. Future work will focus on enabling the model to restructure the DB after iterating over a dataset instead of clustering.
Regarding training data contamination deng-etal-2024-unveiling; samuel-etal-2025-towards, the results of the direct update approach show that the model has not memorised the TOD datasets, as it cannot recall the exact table names for each dialogue.
For larger or structurally different ontologies than TOD, our results show TeQoDO can generalise, though SQL struggles to express multiple hierarchy levels. We aim to scale TeQoDO to diverse ontology structures by adapting the text-to-SQL component; however, its sequential nature limits parallelisation. The large batch size for concatenated Wikipedia articles may affect performance, though evaluating this is computationally expensive. Finally, our approach is sensitive to prompt phrasing, a common issue with large language models razavi2025promptsensitivity.
Our evaluation captures all necessary information and is more fine-grained than existing methods by considering the hierarchy levels of TOD ontologies: domains, slots, values, system actions, and user intents. Using an off-the-shelf similarity model with a fixed threshold may introduce bias, though qualitative analysis is aligned with the quantitative results. In future work, we aim to incorporate human judgment into evaluation to reduce reliance on fixed thresholds.
dennett1987intentional argues that task comprehension is not required for competent performance. Inspired by this, we think that the LLMs’ competence in SQL generation can be used to extract task knowledge in the form of an ontology. We see this work as an important step in distilling human-readable knowledge from text that can be used as a knowledge source for otherwise black-box LLMs. Extracting ontologies automatically with LLMs and analysing their structure may enhance their interpretability in downstream tasks.
@@ 5 Conclusion
We introduce TeQoDO, a method that uses large language models to build ontologies from task-oriented data, exploiting the inherent SQL programming abilities of LLMs and incorporating modular TOD concepts. We evaluate TeQoDO on two widely used TOD datasets and find that it surpasses current SOTA. We also show that TeQoDO generalises to a variety of LLMs and to large ontologies with different structures. Our results motivate further exploration of this approach for explainability.
Acknowledgements
We thank the reviewers and the action editor for their insightful comments and suggested revisions, which significantly improved the quality of this paper. This work was funded by the European Research Council (ERC) under the Horizon 2020 research and innovation program (Grant No. STG2018 804636) and the Ministry of Culture and Science of North Rhine-Westphalia within the Lamarr Fellow Network. Computational resources were provided by the Centre for Information and Media Technology at Heinrich Heine University Düsseldorf and Google Cloud.
@@ A.1 TeQoDO Prompt
    Identify which tables from {db_result_input} are relevant to the new dialogue(s) based on: User intents expressed in the dialogue(s). System actions performed in response. Specific information or entities being queried or discussed. For each relevant table, generate the following SQLite command to inspect its schema: PRAGMA table_info(<table_name>); This will allow you to understand the structure (columns and data types) of the tables you will be working with. Do not create or modify tables yet — only inspect existing ones using PRAGMA. Make sure to wrap all SQL between “‘sql and “‘ for it to be parsed.
@@
Baseline: $\mathcal{C}_{i}=L(d_{i}),\qquad\forall d_{i}\in\mathcal{D},\qquad\mathcal{O}_{i}=\mathcal{C}_{0}\cup\dots\cup\mathcal{C}_{i},\qquad\mathcal{O}_{0}=\{\varnothing\}$
Iterative Baseline: $\mathcal{C}_{i}=L(d_{i},\mathcal{O}_{i-1}),\qquad\mathcal{O}_{i}=\mathcal{O}_{i-1}\cup\mathcal{C}_{i},\qquad\mathcal{O}_{0}=\{\varnothing\}$
Tracking: $\mathcal{C}_{i}=L(d_{i},\mathbf{b}_{i}),\qquad\mathbf{b}_{i}=$ $\mathcal{O}_{i-1}(d_{i})=L(d_{i},\mathcal{O}_{i-1},L(d_{i},\mathcal{O}_{i-1}))$ , $\qquad\mathcal{O}_{i}=\mathcal{O}_{i-1}\cup\mathcal{C}_{i},\qquad\mathcal{O}_{0}=\{\varnothing\}$
Success: $\mathcal{C}_{i}=L(d_{i},\mathbf{b}_{i})$,   $\mathbf{b}_{i}=\mathcal{O}_{i-1}(d_{i},\text{is\_successful}(d_{i})),\qquad\mathcal{O}_{i}=\mathcal{O}_{i-1}\cup\mathcal{C}_{i},\qquad\mathcal{O}_{0}=\{\varnothing\}$
@@ A.2 Predicted Ontology Excerpts
Figure 6: JSON representation of “restaurant” table predicted via TeQoDO on MultiWOZ test set.
@@
Figure 7: JSON representation of “flights” table predicted via TeQoDO on SGD test set.
@@ A.3 Example Dialogues
See Figure˜8 for the example input dialogues from MultiWOZ test set for the generated queries in Table˜6.